OpenStudy (anonymous):

Doug biked to the lake at an average speed of 15 miles per hour. He returned home along the same route at an average speed of 10 miles per hour. If Dough spent a total of 4 hours biking to and from the lake, how many total miles did Dough bike?

OpenStudy (anonymous):

time is distance divide by rate, so time going is \(\frac{D}{15}\) and time returning is \(\frac{D}{10}\) for a total time of \[\frac{D}{15}+\frac{D}{10}\] which you are told is 4 hours solve \[\frac{D}{15}+\frac{D}{10}=4\] for \(D\)